### v1

# Which Demographics do LLMs Default to During Annotation?

Abstract
Demographics and cultural background of annotators influence the labels they assign in text annotation – for instance, an elderly woman might find it offensive to read a message addressed to a “bro”, but a male teenager might find it appropriate. It is therefore important to acknowledge label variations to not underrepresent members of a society. Two research directions developed out of this observation in the context of using large language models (LLM) for data annotations, namely (1) studying biases and inherent knowledge of LLMs and (2) injecting diversity in the output by manipulating the prompt with demographic information. We combine these two strands of research and ask the question to which demographics an LLM resorts to when no demographics is given. To answer this question, we evaluate which attributes of human annotators LLMs inherently mimic. Furthermore, we compare non-demographic conditioned prompts and placebo-conditioned prompts (e.g., “you are an annotator who lives in house number 5”) to demographics-conditioned prompts (“You are a 45 year old man and an expert on politeness annotation. How do you rate {instance}”). We study these questions for politeness and offensiveness annotations on the Popquorn data set, a corpus created in a controlled manner to investigate human label variations based on demographics which has not been used for LLM-based analyses so far. We observe notable influences related to gender, race, and age in demographic prompting, which contrasts with previous studies that found no such effects.
[…]
booktabs
Which Demographics do LLMs Default to During Annotation?
Christopher Bagdon, Aidan Combs, Lynn Greschner, Roman Klinger, Jiahui Li, Sean Papay, Nadine Probol, Yarik Menchaca Resendiz, Johannes Schäfer, Aswathy Velutharambath, Sabine Weber,  and Amelie Wührlthanks: All authors contributed equally. Fundamentals of Natural Language Processing, University of Bamberg {firstname.lastname@uni-bamberg.de}
1 Introduction
In some text annotation tasks, it is feasible to obtain an aggregated ground truth label, for instance in named entity annotation (Yadav and Bethard, 2018) or semantic role labeling (Shi et al., 2020). In other tasks, perhaps even in the majority of tasks, it is more obvious that annotator’s traits influence the label assignments, for instance in sentiment annotation (Liu, 2012), emotion annotation (Klinger, 2023; Plaza-del Arco et al., 2024), or personality profiling (Neuman, 2015). The diversity of annotations, conditioned on annotators’ profiles, has been recognized as an important variable to consider, instead of aggregating all labels to an adjudicated score, which might not correspond to any of the annotators (Plepi et al., 2022).
With the advent of (instruction-tuned) large language models, automatic data annotation and zero or few shot predictions became more popular (Brown et al., 2020). However, language models do not provide the same diversity as human annotators do by simply prompting them in a zero-shot setup, which can lead to a lower performance of the labeling process (Bagdon et al., 2024). To mitigate this problem, it is important to understand whether large language models exhibit a default persona when acting as annotators and how this can be controlled to ensure greater diversity in annotations. One idea to address this limitation is socio-demographic prompting (Muscato et al., 2024), where we specify characteristics like age, gender, ethnicity or socio-economic status.
Previous research has explored the effects of using prompts informed by demographic or cultural contexts on model predictions, but did not find consistent patterns (Beck et al., 2024; Mukherjee et al., 2024). We build on top of this work and particularly contribute in two directions of socio-demographic prompting in a large demographic-annotated data set, namely Popquorn (Pei and Jurgens, 2023) that did not undergo an analysis from the LLM-perspective yet. (1) We study if large language models default to a particular demographic, i.e., we evaluate if the LLM-based predictions default to a particular demographic when not being conditioned. (2) Further, we use the idea of placebos (i.e., irrelevant information that should not change the output of the model) and compare demographic prompts to evaluate their robustness. More concretely, we answer the following questions.
  • RQ1
    What default demographic values can we infer from the annotation behavior of large language models?
  • RQ2
    Are the changes to the models’ annotations more pronounced for demographic prompting than with non-relevant additional information to the prompt?
  • RQ3
    How do task properties of offensiveness rating vs. politeness rating influence the role of demographic information or are observed patterns consistent across the two application tasks?
  • RQ4
    Are observed patterns consistent across different models?
The remainder of the paper is structured as follows: We review related work in Section 2 and explain our data set choices and prompt setup in Section 3. Section 4 presents the results and Section 5 concludes and explains possible future work.
2 Related Work
This paper relates to various areas which we review in the following, namely biases in language models, perspectivism, and (socio-demographic) prompting.
2.1 Inherent Knowledge and Biases in LLMs
Large language models may be understood as databases that store information encoded in the training data (Petroni et al., 2019). These data is stored in a probabilistic manner, which allows these models to show sometimes unexpected generalization capabilities beyond the original instructions available in the training data, i.e., emergent abilities (Wei et al., 2022). It has further been shown that prompting may show properties of adapting parameters through backpropagation purely during inference (von Oswald et al., 2023). Therefore, it is reasonable to consider prompting even as a programming task (Beurer-Kellner et al., 2023).
There are cases in which the information requested from a large language model is not available in the training data “as is” and the generalization abilities reach its limits. In such cases, the model might output text that is not correct. Such cases are sometimes considered “hallucinations”. The confidence of hallucinations is often lower than for correct information (Farquhar et al., 2024).
In cases, in which subjective properties are requested, the models therefore rely on “emergent” abilities. There is some research in understanding the “stances” or “worldviews” encoded in language models. For instance, Motoki et al. (2023); Feng et al. (2023) study political biases and how these lead to unfair models. Ceron et al. (2024) study the reliability and robustness of such stances. Wright et al. (2024) extend such analysis to more fine-grained opinions. The opinions expressed by models correspond more to particular subsets of a society than to others (Santurkar et al., 2023).
2.2 Perspectivism
Natural language processing for a long time focused on seemingly objective tasks such as parsing or named entity recognition. Therefore, it has been a standard approach to adjudicate annotations into a single gold standard (Stubbs, 2011). To do so, a set of methods for aggreggation that considers the distributions and disagreements have been developed (Paun et al., 2018). More recently, research moved towards acknowledging the importance of disagreements. Plank (2022) has been one of the first to discuss challenges and potential approaches to this problem. Recent work looked into methods to consider disagreements (Fleisig et al., 2024) and discusses ethical considerations (Valette, 2024). Nowadays, this field perspectivism developed dedicated shared tasks (Uma et al., 2021) and workshops (Abercrombie et al., 2024).
Disagreement is related to confidence of annotators (and models), an aspect that received some attention. Baumler et al. (2023) made use of this property in the context of active learning, to understand which instances require annotations by multiple people, to understand the disagreements. Troiano et al. (2021); Baan et al. (2024) showed that annotator’s own confidence predicts inter-annotator agreement scores.
To understand label disagreement better, an increasing number of corpora is published with more detailed information on the annotators. Troiano et al. (2023) annotated event reports for emotions and appraisals and collected demographic information, personality, current emotional state of both the person who lived through the event and multiple annotators that read the event description. Plepi et al. (2022) studied the role of demographics, automatically extracted from the data, on the perception of social norms. Bizzoni et al. (2022) discusses the role of individual differences on the judgement of literary quality. Romberg (2022) integrates perspectivism in argument mining, by making explicit the subjective nature of argument interpretation. May et al. (2024) study the role of numbers in social judgements, under consideration of demographics. Frenda et al. (2023) study the perception of irony, depending on nationality, employment status, student status, ethnicity, age, and gender. Sachdeva et al. (2022) measures aspects of hate speech, including sentiment, disrespect, insult, attacking/defending, humiliation, inferior/superior status, dehumanization, violence, genocide, and a 3-valued hate speech benchmark label. They study these variables under the condition of identity target groups and annotator demographics. Xu et al. (2023) look into disagreement of legal case outcome differences. Next to subjective tasks, perspectivism has also been considered in seemingly objective tasks, for instance in named entity recognition (Peng et al., 2024) and natural language inference (Gruber et al., 2024).
The large Popquorn corpus (Pei and Jurgens, 2023) has been created to particularly study perspectivism of annotators and consider the tasks of offensiveness detection, question answering, text rewriting and style transfer, and politeness rating, depending on demographic variables. We use this corpus because it has been created for the study of perspectivism, but it did not undergo analyses from the perspective of large language models.
2.3 Prompting for Automatic Data Set Annotation or Zero-Shot Predictions
While fine-tuning models is currently still the state-of-the-art approach to obtain the best possible performance for a variety of natural language processing tasks, prompting language models for a zero-shot of few-shot prediction gained popularity recently. This is due to the possibility of efficiently adapting model outputs by prompt engineering, without fine-tuning the model – in fact, prompt optimization (by a human or automatically) can be seen as parameter-efficient model adaptation.
This field builds on top of instance-based classification methods and became popular with the work by Palatucci et al. (2009) who suggest to perform dataless classification by semantically encoding output concepts. Another non-NLP example is Banerjee et al. (2022) who embed emotion concepts for zero-shot classification of body gestures. A milestone in the natural language processing community is the work by Yin et al. (2019) who show how natural language inference can be applied across multiple classification tasks and Brown et al. (2020) who show that auto-regressive language models are zero-shot learners.
Since then, a set of studies have been proposed, including work that focuses on cross-linguality (Bareiß et al., 2024), data augmentation (Chen and Shu, 2023), emotions (Plaza-del Arco et al., 2022; Bagdon et al., 2024), named entity recognition (Shen et al., 2023), and sentiment classification Fei et al. (2023); Ma et al. (2022). A more comprehensive survey has been provided by Li (2023). Prompts can also be learned, but this setup is out of scope for our work in this paper (Liu et al., 2023).
2.4 Socio-Demographic Prompting
The automatic annotation with language models does not correspond to human annotation – humans have previous world knowledge, experiences, and perspectives on a matter that differ individually. This comes with issues of automatic annotation. For instance, Bagdon et al. (2024) showed that the diversity of annotations that is beneficial in comparative annotations by multiple people is not straight-forwardly replaced by multiple runs of a language model. Lee et al. (2023) focused on this aspect and studied the alignment of the distribution of labels with human annotataed labels on natural language inference tasks.
Most relevant for our work are the following studies. Beck et al. (2024) study the impact of demographic information on subjective annotation tasks. They find that model variation is larger across other parameters than demographic information (e.g., age and gender) in the prompt. Mukherjee et al. (2024) study a similar setup, namely for cultural aspects of the annotators (e.g., food preferences). To understand if the variables influence the annotation in a systematic way, they compare the predictions to what they call “placebos” – information that should not influence the prediction but looks like relevant parameters (favorite planet or house number). Finally, there is a set of studies that investigate biases on large language models (Cheng et al., 2023; Santy et al., 2023, i.a.).
Our work combines aspects of previous research, namely on biases, placebos, and demographics.
3 Experimental Setting
This section presents the methodology and resources utilized in our experiments.
3.1 Data Sets
We chose the Popquorn data for our experiments (Pei and Jurgens, 2023). In contrast to data used by other previous research we are aware of, these data has been particularly sampled for the study of annotator’s properties and their impact on annotation tasks. Therefore, these data render themselves as a straight-forward choice also for an LLM-based analysis. The original data stems from the Ruddit corpus (Hada et al., 2021) which has originally been annotated via best–worst scaling. BWS may lead to different label frequencies than rating scale annotations and is not appropriate for skewed distributions (Finn and Louviere, 1992; Louviere et al., 2015). The authors of Popquorn reannotated with 1–5 rating scales, which corresponds to our annotation setup. Further, it is noteworthy that the creators of the data set removed annotators with a particularly low annotation competency according to Mace (Hovy et al., 2013), to filter out potential random noise.
The Popquorn data set consists of 45,000 annotations from 1,484 annotators and includes information on the annotators on gender, race, age, occupation and education. It is partioned into multiple NLP tasks, from which we use the task on offensiveness ratings and politeness ratings. To equalize the number of annotations per instance, we sample three annotations for each instance. When sampling, we omit the annotations where the annotator preferred not to disclose demographic information. We exclude a data instance because it has fewer than three annotations remaining. The two subsets of the sample which we use in our experiments consist of 4,500 annotations of 1,500 instances for offensiveness ratings and 11,151 annotations of 3,717 instances for politeness ratings.
3.2 Prompt Setup
We use three different prompt setups, namely socio-demographically informed (SD) prompts, placebo-informed (P) prompts, and prompts without additional information (N prompts). We show the templates for these prompts in Table 1. The input to the LLM consist of a concatenation of the system message with the respective text to be classified. The values for demographic attributes are sampled from the demographics of the annotators. The sets of values for the placebo attributes are displayed in Table 2. During prompting, each placebo attribute value is randomly sampled from the respective set.
Prompt Type Prompt 1 Template Prompt 2 Template
Socio-Demographic (SD) You are a person of gender [gender], race [race], age [age], occupation [occupation], education [education]. Given a text, how do you rate the [task] of the following text? Possible values are [task-values]. In your answer, only provide the label you chose. Imagine [Prompt 1 with initial character in lower case]
Placebo (P) You are a person of height [height], Zodiac sign [zodiac sign], living in a house with the number [house number], enjoying [enjoying], and whose favorite color is [favorite colour]. Given a text, how do you rate the [task] of the following text? Possible values are [task-values]. In your answer, only provide the label you choose. Imagine [Prompt 1 with initial character in lower case]
Non-Demographic (N) You are a person of any gender, any race, any age, with any occupation, and any education. Given a text, how do you rate the [task] of the following text? Possible values are [task-values]. In your answer, only provide the label you choose. Given a text, how do you rate the [task] of the following text? Possible values are [task-values]. In your answer, only provide the label you chose.
Table 1: Prompt types and Templates for system messages. For the [task] offensiveness, the [task-values] are ‘not offensive’, ‘slightly offensive’, ‘moderately offensive’, ‘very offensive’ or ‘extremely offensive’. For the [task] politeness, the [task-values] are ‘not polite’, ‘slightly polite’, ‘moderately polite’, ‘very polite’ or ‘extremely polite’.
Placebo Attribute Possible Values
Height 140 cm, 150 cm, 160 cm, 170 cm, 180 cm, 190 cm, 200 cm, 210 cm.
Zodiac sign Aries, Taurus, Gemini, Cancer, Leo, Virgo, Libra, Scorpio, Sagittarius, Capricorn, Aquarius, Pisces.
House number 6, 12, 13, 24, 45, 68, 98, 122, 234, 1265.
Enjoying food, sleep, friends.
Favorite colour red, green, blue, yellow, purple, turquoise, orange, pink, black, white, brown.
Table 2: Values of placebo attributes used in our P prompts.
3.3 Model Choice and Access
We use two LLMs, namely GPT-4o [^1] https://openai.com/index/hello-gpt-4o/ and Claude [^2] https://www.anthropic.com/news/claude-3-5-sonnet, claude-3-5-sonnet-20240620. We access both models via their respective APIs. The total cost for using GPT-4o amounted to \$54, while the cost for using Claude was \$60.
3.4 LLM Output Parsing
We transform the texts generated by the LLMs into a defined label set of five categories, corresponding to a rating scale of 1 to 5 for each of the two tasks. We used the Langchain StrOutputParser [^3] https://api.python.langchain.com/en/latest/output_parsers/langchain_core.output_parsers.string.StrOutputParser.html to interpret the output from GPT-4o as one of the designated labels, while the output from Claude was utilized directly.
There were few cases where the output could not be parsed. We encountered 22 such cases with GPT-4o and three with Claude. Of the failed cases with GPT-4o, 14 were due to the input text being in Polish instead of English. This resulted in the model predicting labels in Polish, despite the prompt specifying an English-only label set. The remaining eight cases involved very short text samples or questions, which the model misidentified and prompted it to request clarification instead of performing the classification task.
For Claude, the three instances of failure involved actor names in the text. This triggered the model’s copyright protection protocols, preventing it from reproducing or paraphrasing potentially copyrighted content. Given the total number of instances analyzed in our experiments, these numbers are relatively insignificant. Consequently, we disregarded the output in these cases and assigned the labels “not offensive” or “not polite”.
3.5 Evaluation Settings
We conduct various analyses of the labels generated by the LLMs, utilizing the different prompt setups described in Section 3.2. First, we evaluate the extent to which the LLMs replicate human annotations by comparing the predictions made with non-demographic (N) prompts to the gold standard annotations. Second, we examine the impact of socio-demographic prompting by comparing the models’ automatic annotations produced with SD prompts to those produced with N prompts. In this analysis, we assess the differences in the LLMs’ annotations when demographic data is included versus when it is omitted. Thirdly, we investigate the effect of placebo prompting in a similar manner by comparing the automatic annotations generated with P prompts to those created with N prompts.
4 Results
In this section, we present the results of our experiments based on the annotations generated by the LLMs for two labeling tasks. Each of our research questions is addressed with particular results derived from the various experiments conducted. Detailed results for individual prompts given in Section A.2 show that the two corresponding prompts for no-info promping behave similarly. We also observed this for the other prompt types (see Table 1). Thus, we here report all results as an average over the two respective prompts for the different prompt types.
Offensiveness Politeness
Socio-Demographic Attribute GPT-4o Claude GPT-4o Claude
Age **0.01 **0.01 0.00 0.00
Gender (ref.: Male)
Female 0.00 --0.03 --0.05 --0.05
Non-binary --0.06 --0.01 --0.05 --0.05
Race (ref.: White)
Asian 0.09 0.03 --0.08 0.00
Black/Afri. Am. ***0.22 **0.19 **0.14 **0.15
Hispanic or Latino --0.11 --0.05 0.09 0.12
Other race --0.14 --0.26 --0.17 --0.15
Occupation (ref.: Employed)
Unemployed 0.04 --0.08 --0.08 --0.08
Homemaker --0.07 --0.02 --0.03 --0.06
Retired --0.11 --0.13 0.03 0.03
Self-employed 0.04 0.02 --0.03 --0.03
Student 0.13 0.13 --0.11 --0.13
Other occupation --0.05 0.02 0.08 0.11
Education (ref.: Less than high school)
High school dipl. --0.01 --0.08 --0.34 *--0.37
College degree 0.05 --0.09 *--0.43 *--0.48
Graduate degree 0.06 --0.01 *--0.36 *--0.44
Other education --0.02 0.01 *--0.50 **--0.57
Table 3: Coefficients indicating the effect of particular human demographic categories on the distance between human and LLM annotations. Statistical significance is calculated based on standard error (see Section A.1) and is here marked by asterisks: * corresponds to P $\leq$ 0.05, ** to P $\leq$ 0.01, and *** to P $\leq$ 0.001.
4.1 RQ1: What default demographic values can we infer from the annotation behavior of large language models?
We approach the identification of the default persona of models in two different ways following the first two settings as described in Section 3.5. The following sections present and discuss the results of these analyses.
Which socio-demographic attributes of human annotators does an LLM inherently mimic in the absence of explicit information (N prompts)? Table 3 shows the results for this analysis which corresponds to the first setting described in Section 3.5. We combine socio-demographic attributes that are represented in only a few cases. The table specifies the reference categories (ref.) used for each of the categorical variables. We define such reference to report coefficients that describe the relation to other socio-demographic categories.
Table 3 shows coefficients for the two models (GPT-4o and Claude) and two classification tasks. The coefficients can be interpreted as the effect of particular human demographic categories on the distance between human and LLM annotations. The dependent variables to compute the scores are the difference between the absolute values of the human annotation and the annotation of the respective LLM with non-demographic prompting. The independent variables are the human annotator demographic attribute values. Positive coefficients indicate that the LLM is further from the human annotators in that demographic category, i.e., less accurate. So, categories with positive, statistically significant coefficients are those categories which the respective model does not default to.
This analysis (see Table 3) reveals significant biases in the LLMs’ predictions based on demographic factors, particularly concerning race and age. Specifically, the models demonstrate a tendency to align more closely with annotations provided by persons identying as White as opposed to those with the value Black or African American. The distance between their annotations measures .19 to .22 Likert scale points for offensiveness rating and .14 to .15 points for politeness rating. Additionally, the LLMs are progressively less accurate in reflecting the views of older individuals at offensiveness rating, with a .01 point increase in distance for each year of age. In terms of educational background, the models exhibit a greater discrepancy from those with less than a high school education compared to those with higher educational levels, ranging from .34 to .57 points for politeness rating. These patterns suggest that the models’ training data may lack representation for these demographic groups, affecting its annotation accuracy.
Notably, the analysis found no significant effects related to gender or occupational status. Overall, while the statistically significant differences in annotation distances suggest some demographic biases, these differences are modest compared to the full Likert scale range. This complexity makes it challenging to establish a clear default persona for the LLMs. Therefore, we perform a second analysis in the following.
Offensiveness Politeness
Socio-Demographic Attribute count $\Delta_{\mu}$ (GPT-4o) $\Delta_{\mu}$ (Claude) count $\Delta_{\mu}$ (GPT-4o) $\Delta_{\mu}$ (Claude)
Gender
Male 2,157 0.18 0.17 5,195 0.26 0.20
Female 2,219 0.20 0.15 5,623 0.24 0.22
Non-binary 124 0.29 0.17 333 0.24 0.17
Race
White 3,396 0.18 0.16 8,163 0.25 0.20
Hispanic or Latino 95 0.21 0.11 790 0.22 0.21
Native American 97 0.25 0.12 0
Arab American 17 0.32 0.06 0
Native Hawaiian or Pacific Islander 0 36 0.31 0.15
American Indian or Alaska Native 0 28 0.18 0.23
Black or African American 559 0.22 0.15 1,386 0.25 0.21
Asian 336 0.21 0.15 731 0.24 0.24
Hebrew 0 17 0.18 0.26
Age
18-24 499 0.21 0.16 1,241 0.25 0.16
25-29 444 0.20 0.16 894 0.25 0.18
30-34 540 0.16 0.15 1,190 0.24 0.17
35-39 532 0.20 0.19 834 0.26 0.18
40-44 418 0.19 0.18 1,176 0.23 0.21
45-49 388 0.19 0.15 957 0.22 0.22
50-54 314 0.17 0.17 995 0.27 0.21
54-59 627 0.18 0.15 1,083 0.26 0.22
60-64 251 0.22 0.15 1,193 0.24 0.22
>65 487 0.20 0.15 1,588 0.25 0.26
Occupation
Unemployed 571 0.19 0.14 1,328 0.28 0.19
Employed 2,189 0.19 0.17 4,944 0.24 0.20
Homemaker 199 0.24 0.14 784 0.23 0.21
Retired 500 0.19 0.15 1,783 0.26 0.25
Other 86 0.16 0.22 268 0.28 0.19
Self-employed 617 0.17 0.17 1,395 0.23 0.21
Student 338 0.23 0.14 649 0.23 0.16
Education
Less than a high school diploma 84 0.18 0.17 76 0.30 0.22
High school diploma or equivalent 1,379 0.19 0.14 3,312 0.26 0.17
Graduate degree 846 0.21 0.16 2,160 0.23 0.25
College degree 2,098 0.18 0.17 5,352 0.24 0.21
Other 93 0.25 0.11 251 0.24 0.23
Table 4: Sample sizes and mean distance scores of demographic-prompting predictions in comparison to no-info-prompting predictions for models GPT-4o and Claude at two rating tasks. Variables with very few cases are shown in gray to indicate a lack or reliability of these numbers.
The Effect of Demographic Prompting. In a second approach, we examine how the inclusion of demographic information in the prompt (cf. second setting in Section 3.5) influences the automatic annotation outputs of the models. We compare SD prompts with N prompts. This allows us to infer the demographics that the models defaults to.
Table 4 shows the results as differences in the prediction scores ($\Delta_{\mu}$). [^4] Note that there is an overlap in the age ranges (50–54, 54–59), a consequence of the survey question design in the original annotation task for the Popquorn dataset. We have preserved these original categories, despite this issue, in order to maintain the fine-grained distinctions they provide. For some demographic attributes, there are relatively few samples (“count” columns). This implies that there may not be sufficient statistical evidence to support the observed average distance values for certain categories. We focus our discussion on cases with $\geq$100 cases.
GPT-4o shows substantial differences in the scores for the offensiveness task across different gender attribute values: the predictions for the attribute value Non-Binary is substantially larger (.29) than for Male (.18) and Female (.2). This suggests that prompting the LLM to act as a non-binary individual has a more pronounced effect on its predictions, indicating that the male or female attributes are more aligned with its default persona. This observation is not consistent across tasks.
Claude shows a notable pattern for the politeness task concerning the age socio-demographic attribute. With an increasing age, the prediction difference increases (from 16 to .26). Similarly, the occupation attribute reflects the highest difference for the value Retired. This indicates that when Claude is prompted to act as an older individual, it exhibits an increased sensitivity to politeness.
4.2 RQ2: Are the changes to the models’ annotations more pronounced for demographic prompting than with non-relevant additional information to the prompt?
We investigate whether the modifications to the model are more substantial for socio-demographic prompting compared to non-relevant additional information provided in the prompts. Specifically, we analyze how the models’ prediction scores for the two application tasks are affected when presented with placebo information (cf. third setting as given in Section 3.5).
We evaluate if these results are as pronounced as the observations based on Table 4 as described above. Overall, the results of the placebo prompting (P prompts) in comparison to (N prompts) do not reveal any notable differences for specific attribute values. The scores remain consistently stable across these comparisons. Therefore, we show the results only in the appendix, in Section A.3. Consequently, we conclude that the changes in model behavior are indeed more pronounced for socio-demographic prompting compared to the inclusion of non-relevant additional information.
4.3 RQ3: How do task properties of offensiveness rating vs. politeness rating influence the role of demographic information or are observed patterns consistent across the two application tasks?
We investigate how task properties influence the role of demographic information in prompting and whether these patterns remain consistent across different application tasks. Our general conclusion indicates that there is no clearly distinguished default persona of the models for both tasks. Table 3 suggest that the results are in general consistent for the two tasks. However, we did also highlight some task-specific tendencies above at the analysis regarding RQ1. Notably, there are also differences in the absolute values of the prediction distances when comparing demographic prompting to no-information prompting across the two tasks. The average prediction difference associated with socio-demographic prompting for offensiveness rating is calculated to be 0.19. In contrast, the average for politeness rating is substantially higher at .25. Similar results are evident for the model Claude (0.16 for offensiveness rating and 0.21 for politeness rating). This suggests that, in general, the LLMs are more influenced by demographic information at rating politeness than rating offensiveness.
4.4 RQ4: Are observed patterns consistent across different models?
The patterns identified in the analysis presented in Table 3 demonstrate consistency across the two models examined, particularly in their ability to replicate human annotations. Notably, Table 4 highlighted specific behaviors of the LLMs in response to socio-demographic prompting. However, each of these distinct behaviors was only observed in one of the models. Overall, both LLMs tested in this study did not display clear default personas. Thus, the models remain consistent in this aspect.
5 Conclusion and Future Work
With this paper, we presented the first analysis of the effect of socio-demographic prompting on tasks in the Popquorn data set. This corpus has been developed particularly for the study of the impact of demographic variables and therefore also offers itself for LLM-based studies.
We find that demographic prompting has some measurable effects on the annotation behaviors of large language models, contrasting with placebo prompting, which elicited no consistent changes across various attributes. Specifically, our analyses show that LLMs exhibit variations in annotation based on demographic attributes, particularly highlighting influences related to gender, race and age. While we cannot infer one concrete, unique default persona, we conclude that large language models do not represent all members of a society alike, and that socio-demographic prompting does influence the result in a structured manner. This result is in contrast to the findings by previous work that studied other corpora and other models.
The findings are not overall consistent across models. While this could be considered a principled issue, it also presents itself with a substantial challenge: depending on the model that is used by an end-user, the impact of socio-demographic information varies, and different models default to different demographic information.
We therefore propose that future work focuses on the development of reliable instruments to measure the influence of implicitly or explicitly provided information about the user. We presume that making such information explicit and ensuring that it is interpreted in a reliable manner by models is an appropriate approach to mitigate issues with unknown biases that are not transparent to users or automatic processing pipelines.
Limitations
The goal of our work is to develop an overall understanding of the impact of socio-demographic prompting on language models. Naturally, such study can only focus on a limited set of models and variables. Particularly limiting is the set of attributes available in data sets, which are also culturally biases towards the USA. Therefore it is important to conduct more such similar studies across languages, cultures, and regions. Also in our study, a considerable limitation is the unequal representation of demographic subgroups. This disparity can lead to unreliable results and may impact the conclusions drawn from our analyses, particularly regarding the influence of demographic prompting.
We also limited our experiments to only two large language models, conducting each experiment just once, for pragmatic reasons stemming from limited compute resources. A large scale analysis across a multitude of tasks would mitigate such issue.
Finally, it is important to note that the reported differences in annotation outputs are relatively small compared to the overall Likert scale range of 1 to 5. This raises questions about the practical significance of the observed effects, necessitating further exploration to understand their impact in real-world applications.
Ethical Considerations
Our work has the goal to make challenges transparent which the use of large language models has. A considerable limitation from an ethical perspective is that the variables we consider is not relevant across cultures in the world. Particularly the “race” variable is heavily biased towards the US American society. Other variables might be relevant in other cultures in the world that we are not aware of. We therefore suggest to conduct future studies that consider more open sets of variables that describe the users of language model-based systems.
Appendix A Additional Results
In this section we provide more detailed results of our experiments.
A.1 Full Results for LLMs Mimicing Annotators
Offensiveness Politeness
Socio-Demographic Attribute GPT-4 Claude GPT-4 Claude
Intercept 0.45 (0.17)** 0.56 (0.19)** 1.45 (0.19)*** 1.53 (0.20)***
Age (years) 0.01 (0.00)** 0.01 (0.00)** 0.00 (0.00) 0.00 (0.00)
Gender (ref: Male)
Female 0.00 (0.04) -0.03 (0.04) -0.05 (0.03) -0.05 (0.03)
Non-binary -0.06 (0.12) -0.01 (0.13) -0.05 (0.10) -0.05 (0.10)
Race (ref: White)
Asian 0.09 (0.08) 0.03 (0.08) -0.08 (0.07) 0.00 (0.07)
Black or African American 0.22 (0.06)*** 0.19 (0.06)** 0.14 (0.05)** 0.15 (0.05)**
Hispanic or Latino -0.11 (0.14) -0.05 (0.15) 0.09 (0.06) 0.12 (0.06)+
Other race/ethnicity -0.14 (0.13) -0.26 (0.14)+ -0.17 (0.18) -0.15 (0.19)
Occupation (ref: Employed)
Unemployed 0.04 (0.07) -0.08 (0.07) -0.08 (0.05) -0.08 (0.06)
Homemaker -0.07 (0.10) -0.02 (0.10) -0.03 (0.07) -0.06 (0.07)
Retired -0.11 (0.07) -0.13 (0.08) 0.03 (0.06) 0.03 (0.06)
Self-employed 0.04 (0.06) 0.02 (0.07) -0.03 (0.05) -0.03 (0.05)
Student 0.13 (0.08) 0.13 (0.09) -0.11 (0.08) -0.13 (0.08)
Other occupation -0.05 (0.15) 0.02 (0.16) 0.08 (0.11) 0.11 (0.11)
Education (ref: Less than high school)
High school diploma or equivalent -0.01 (0.15) -0.08 (0.17) -0.34 (0.18)+ -0.37 (0.19)*
College degree 0.05 (0.15) -0.09 (0.17) -0.43 (0.18)* -0.48 (0.19)*
Graduate degree 0.06 (0.16) -0.01 (0.17) -0.36 (0.18)* -0.44 (0.19)*
Other education -0.02 (0.21) 0.01 (0.22) -0.50 (0.21)* -0.57 (0.22)**
SD (instance intercepts) 0.44 0.42 0.37 0.33
SD (annotator intercepts) 0.23 0.27 0.30 0.32
Num. Obs. 4500 4481 11151 11151
R2 Marg. 0.014 0.016 0.012 0.014
R2 Cond. 0.316 0.321 0.311 0.305
BIC 11712.2 11495.0 27644.5 27046.1
ICC 0.3 0.3 0.3 0.3
RMSE 0.67 0.66 0.66 0.65
Table 5: Detailed analyses and coefficients indicating the effect of particular human demographic categories on the distance between human and LLM annotations.
Table 5 provides additional statistics complementing those presented in Table 3. Here, we describe these additional statistics in more detail. At the bottom of Table 5, we present regression model fit statistics. The numbers in parentheses represent the standard error of the estimates, which indicates the uncertainty associated with these estimates and is used for calculating statistical significance.
“Intercept” reflects the overall intercept of the model, representing the expected distance between the LLM’s predictions and those of a human, assuming all reference categories and an average age of 0 with an average annotation skill on a prompt of average ambiguity. The other coefficients serve as adjustments to this baseline value. The statistical significance of the intercept itself is not meaningful, and it is the effect sizes that are of primary interest rather than the absolute values of the predicted distances. A “+” sign indicates marginal significance (0.05 $\leq$ P $\leq$ 0.1). SD (Standard Deviation) refers to the standard deviations of the mixed-effect model’s random intercepts for instances and annotators. The model assigns a unique intercept to each instance and annotator to account for uninteresting idiosyncrasies, with a mean of 0. The SD value indicates the expected variability of LLM-human differences across instances and annotators. “Num. Obs.” denotes the number of rows (annotations) on which the regression model was conducted. R2 Marg. and R2 Cond. are measures of the model’s explanatory power. R2 Marginal indicates how much variation is accounted for by the coefficients alone, which is very low, suggesting that these coefficients contribute little to predictive power despite some being statistically significant. R2 Conditional, on the other hand, indicates the proportion of variance explained by both the coefficients and the random intercepts, which is significantly higher, demonstrating that the random intercepts contribute more substantially to the model’s explanatory power.
BIC (Bayesian Information Criterion) is a model fit statistic useful for model comparison, though its value is not meaningful in isolation. It is included to allow for calculations of BIC values for both base and limited models. The ICC (Intraclass Correlation Coefficient) indicates the extent to which the clustering (by instances and annotators) influences the outcomes (distances between human and LLM ratings). It represents the ratio of total variance in the dependent variable attributed to variance between cluster means, as opposed to variance within clusters. A value of 0.3 or higher suggests that the inclusion of random intercepts is necessary to account for clustering. If the ICC were very low (less than 0.1), a simpler model could be justified.
Finally, RMSE (Root Mean Square Error) represents the standard deviation of the residuals; however, it provides limited utility in isolation.
A.2 Results for Individual Prompts
In examining the scores for the individual prompts presented in Table 6 and Table 7, we observe that there are no notable differences between the two prompts. This lack of variation in scores is a positive outcome, as it suggests consistency in the model’s performance across different prompts. Such uniformity reinforces the reliability of the results, indicating that the prompts used do not unduly influence the models’ annotation behaviors. This consistency allows us to have greater confidence in our findings and their implications regarding the impact of demographic information on model outputs.
Offensiveness Prompt 1 Offensiveness Prompt 2
Socio-Demographic Attribute GPT-4 Claude GPT-4 Claude
Intercept 0.51 (0.17)** 0.60 (0.18)*** 0.39 (0.18)* 0.52 (0.20)**
Age (years) 0.01 (0.00)** 0.00 (0.00)** 0.01 (0.00)** 0.01 (0.00)***
Gender (ref: Male)
Female 0.00 (0.04) -0.03 (0.04) 0.00 (0.04) -0.02 (0.05)
Non-binary -0.08 (0.12) 0.01 (0.13) -0.05 (0.13) -0.02 (0.14)
Race (ref: White)
Asian 0.12 (0.07) 0.03 (0.08) 0.06 (0.08) 0.02 (0.09)
Black or African American 0.21 (0.06)*** 0.19 (0.06)** 0.22 (0.06)*** 0.19 (0.07)**
Hispanic or Latino -0.08 (0.14) -0.03 (0.15) -0.15 (0.15) -0.07 (0.16)
Other race -0.11 (0.13) -0.23 (0.14)+ -0.18 (0.13) -0.29 (0.15)*
Occupation (ref: Employed)
Unemployed 0.03 (0.07) -0.08 (0.07) 0.05 (0.07) -0.07 (0.08)
Homemaker -0.08 (0.10) 0.01 (0.10) -0.05 (0.10) -0.05 (0.11)
Retired -0.11 (0.07) -0.12 (0.08) -0.11 (0.08) -0.14 (0.09)+
Self-employed 0.02 (0.06) 0.03 (0.06) 0.05 (0.06) 0.01 (0.07)
Student 0.11 (0.08) 0.11 (0.09) 0.15 (0.09)+ 0.16 (0.09)+
Other occupation -0.07 (0.14) 0.04 (0.15) -0.04 (0.15) 0.01 (0.16)
Education (ref: Less than high school)
High school diploma or equivalent -0.04 (0.15) -0.08 (0.16) 0.02 (0.16) -0.07 (0.17)
College degree 0.03 (0.15) -0.09 (0.16) 0.07 (0.16) -0.09 (0.18)
Graduate degree 0.02 (0.16) -0.01 (0.17) 0.10 (0.17) -0.01 (0.18)
Other education -0.03 (0.20) 0.04 (0.22) 0.00 (0.22) -0.03 (0.23)
SD (instance intercepts) 0.48 0.43 0.43 0.41
SD (annotator intercepts) 0.22 0.26 0.25 0.29
Num.Obs. 4500 4481 4500 4481
R2 Marg. 0.012 0.014 0.015 0.018
R2 Cond. 0.325 0.314 0.301 0.317
BIC 12097.9 11684.3 11899.0 11699.0
ICC 0.3 0.3 0.3 0.3
RMSE 0.69 0.68 0.70 0.68
Table 6: Results of the analysis of individual prompts for the offensiveness rating task. This table displays coefficients for the two models (GPT-4o and Claude) and indicates the effects of specific socio-demographic attributes of human annotators on the discrepancies between human and LLM annotations. Positive coefficients signify that the LLM is less accurate in mimicking the responses of annotators from certain demographic categories.
Politeness Prompt 1 Politeness Prompt 2
Socio-Demographic Attribute GPT-4 Claude GPT-4 Claude
Intercept 1.40 (0.19)*** 1.43 (0.19)*** 1.51 (0.19)*** 1.65 (0.21)***
Age (years) 0.00 (0.00) 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Gender (ref: Male)
Female -0.05 (0.03) -0.05 (0.03) -0.05 (0.03) -0.05 (0.04)
Non-binary -0.07 (0.10) -0.05 (0.10) -0.04 (0.10) -0.05 (0.11)
Race (ref: White)
Asian -0.08 (0.06) 0.00 (0.07) -0.08 (0.07) -0.02 (0.07)
Black or African American 0.13 (0.05)** 0.15 (0.05)** 0.15 (0.05)** 0.16 (0.05)**
Hispanic or Latino 0.07 (0.06) 0.14 (0.06)* 0.10 (0.06)+ 0.11 (0.07)
Other race -0.23 (0.18) -0.15 (0.18) -0.11 (0.18) -0.15 (0.20)
Occupation (ref: Employed)
Unemployed -0.08 (0.05) -0.07 (0.05) -0.07 (0.05) -0.10 (0.06)+
Homemaker -0.03 (0.07) -0.05 (0.07) -0.03 (0.07) -0.06 (0.07)
Retired 0.04 (0.06) 0.04 (0.06) 0.02 (0.06) 0.03 (0.06)
Self-employed -0.02 (0.05) -0.02 (0.05) -0.04 (0.05) -0.03 (0.06)
Student -0.10 (0.08) -0.11 (0.08) -0.11 (0.08) -0.14 (0.08)+
Other occupation 0.08 (0.11) 0.10 (0.11) 0.08 (0.11) 0.11 (0.12)
Education (ref: Less than high school)
High school diploma or equivalent -0.31 (0.18)+ -0.35 (0.18)+ -0.37 (0.18)* -0.40 (0.20)*
College degree -0.41 (0.18)* -0.45 (0.18)* -0.46 (0.18)* -0.52 (0.20)**
Graduate degree -0.33 (0.18)+ -0.41 (0.18)* -0.40 (0.19)* -0.49 (0.20)*
Other education -0.46 (0.21)* -0.56 (0.21)** -0.54 (0.21)* -0.60 (0.23)**
SD (instance intercepts) 0.41 0.34 0.44 0.40
SD (annotator intercepts) 0.30 0.30 0.30 0.34
Num.Obs. 11151 11151 11151 11151
R2 Marg. 0.011 0.013 0.010 0.013
R2 Cond. 0.322 0.292 0.340 0.340
BIC 28301.7 27294.4 28833.6 28424.2
ICC 0.3 0.3 0.3 0.3
RMSE 0.67 0.66 0.67 0.67
Table 7: Results of the analysis of individual prompts for the politeness rating task. This table presents the coefficients for the two models (GPT-4o and Claude), demonstrating how different socio-demographic characteristics of human annotators influence the distance between human and LLM annotations. Positive coefficients indicate a lower accuracy of the LLM in reflecting the views of annotators from particular demographic categories.
A.3 Results for Placebo Prompting
Table 8 presents the results for placebo prompting. It includes sample sizes and mean distance scores ($\Delta_{\mu}$) for predictions generated using placebo prompts compared to those produced with no-info prompts for the models GPT-4o and Claude across two rating tasks. The results show that placebo prompting (P prompts) does not yield any notable differences relative to no-info prompting (N prompts) for specific attribute values. Overall, the scores remain stable across these comparisons. This consistency suggests that introducing non-relevant information does not change the models’ performance systematically. It reinforces our conclusion that the notable changes in model behavior are primarily driven by demographic prompting rather than by placebo or non-relevant prompts.
Offensiveness Politeness
Placebo Attributes count $\Delta_{\mu}$ (GPT-4o) $\Delta_{\mu}$ (Claude) count $\Delta_{\mu}$ (GPT-4o) $\Delta_{\mu}$ (Claude)
Height
140 cm 601 0.23 0.19 1380 0.27 0.19
150 cm 555 0.27 0.20 1393 0.27 0.18
160 cm 579 0.23 0.19 1417 0.26 0.20
170 cm 536 0.20 0.20 1407 0.27 0.18
180 cm 574 0.24 0.20 1354 0.24 0.18
190 cm 564 0.28 0.23 1385 0.25 0.18
200 cm 572 0.24 0.24 1446 0.27 0.18
210 cm 519 0.25 0.22 1369 0.25 0.18
Zodiac sign
Aries 395 0.24 0.22 889 0.25 0.17
Taurus 382 0.26 0.20 913 0.27 0.18
Gemini 355 0.21 0.23 895 0.26 0.18
Cancer 383 0.29 0.27 956 0.27 0.18
Leo 378 0.24 0.22 928 0.27 0.19
Virgo 360 0.22 0.23 915 0.26 0.19
Libra 398 0.23 0.19 974 0.24 0.19
Scorpio 390 0.25 0.18 975 0.26 0.19
Sagittarius 352 0.23 0.20 919 0.27 0.19
Capricorn 387 0.25 0.21 983 0.27 0.18
Aquarius 365 0.28 0.21 892 0.25 0.17
Pisces 355 0.23 0.17 912 0.24 0.20
House number
6 460 0.26 0.21 1130 0.28 0.19
12 446 0.24 0.22 1090 0.26 0.18
13 447 0.23 0.21 1107 0.26 0.19
24 424 0.27 0.22 1103 0.25 0.18
45 455 0.24 0.18 1123 0.26 0.19
68 438 0.22 0.19 1098 0.26 0.18
98 456 0.27 0.23 1190 0.25 0.20
122 465 0.23 0.21 1116 0.26 0.19
234 466 0.23 0.21 1118 0.26 0.19
1265 443 0.26 0.22 1076 0.26 0.17
Enjoying
food 1468 0.25 0.22 3793 0.26 0.19
sleep 1486 0.23 0.21 3662 0.26 0.18
friends 1546 0.25 0.21 3696 0.26 0.19
Favorite colour
red 436 0.25 0.22 1046 0.28 0.16
green 399 0.27 0.24 1013 0.26 0.18
blue 400 0.23 0.21 997 0.27 0.19
yellow 444 0.24 0.22 984 0.25 0.19
purple 443 0.27 0.20 1055 0.25 0.20
turquoise 367 0.23 0.20 1056 0.23 0.21
orange 405 0.23 0.21 1033 0.26 0.18
pink 371 0.25 0.21 976 0.28 0.16
black 441 0.21 0.20 993 0.25 0.17
white 381 0.27 0.21 991 0.28 0.20
brown 413 0.21 0.21 1007 0.26 0.19
Table 8: Sample sizes and mean distance scores of placebo prompting predictions in comparison to no-info prompting predictions for models GPT-4o and Claude at two rating tasks.